5. a) Write the expression x1 ∨x2 ∧x3 ∨x4 in conjunction normal form and disjunctive

normal form


1
Expert's answer
2021-04-15T07:25:23-0400

Let's write the expression in conjunctive normal form:

"{x_1} \vee {x_2} \wedge {x_3} \vee {x_4} = {x_1} \vee {x_4} \vee {x_2} \wedge {x_3} = \left( {{x_1} \vee {x_2} \vee {x_4}} \right) \wedge \left( {{x_1} \vee {x_3} \vee {x_4}} \right)"

"{x_1},\,{x_2} \wedge {x_3},\,\,{x_4}\," are elementary conjunctions, so the expression is already written in disjunctive normal form: "{x_1} \vee {x_2} \wedge {x_3} \vee {x_4} = {x_1} \vee \left( {{x_2} \wedge {x_3}} \right) \vee {x_4}"

Answer: CNF: "\left( {{x_1} \vee {x_2} \vee {x_4}} \right) \wedge \left( {{x_1} \vee {x_3} \vee {x_4}} \right)" , DNF: "{x_1} \vee \left( {{x_2} \wedge {x_3}} \right) \vee {x_4}"